A bicycle of mass 14.6 kg and a rider of mass 50 kg generate a net force of 12 N. how fast are they going after 2.0s?

1
Expert's answer
2022-04-12T16:12:51-0400

The acceleration of a bicycle

"a=\frac{F}{m}=\frac{12}{50+14.6}=0.186\:\rm m/s^2"

The velocity of the bicycle after 2.0 s

"v=at\\
=0.186*2.0=0.37\:\rm m/s=1.3\: km/h"